OpenStudy (anonymous):

3 questions i will give MEDAL and FAN if you give me the answers !!!!!

OpenStudy (anonymous):

Find the value of x. Round to the nearest degree.

OpenStudy (anonymous):

Write the ratios for sin X and cos X.

OpenStudy (anonymous):

Describe the vector as an ordered pair. Round the coordinates to the nearest tenth

OpenStudy (anonymous):

1. sin(x) = 7/9 sin^-1(7/9) = x x = 51 degrees 2. sin(X) = sqrt(119)/12 cos(X) = 5/12 3. (19.4,50.4)

OpenStudy (anonymous):

thank you <3